OpenStudy (anonymous):

If the reaction of 4.50 moles of sodium with excess hydrofluoric acid produced an 85.0% yield of hydrogen gas, what was the actual yield of hydrogen gas? Unbalanced equation: Na + HF “yields”/ NaF + H2 1.91 mol H2 2.25 mol H2 2.65 mol H2 3.82 mol H2

OpenStudy (aaronq):

find the theoretical yield, multiply it by 85%.

OpenStudy (anonymous):

can you help me find the theoretical yield?

OpenStudy (aaronq):

First you need to balance the equation. Then use a ratio of the moles of the limiting reagent (sodium in this case) and the moles of product with their stoichiometric coefficients. \(\sf \dfrac{moles~of~Na}{Na's~coefficient}=\dfrac{moles~ofH_2}{H_2's ~coefficient}\)

OpenStudy (anonymous):

Because I got 4.536 and I am not sure if that is the theoretical yield.

OpenStudy (aaronq):

that's not right

OpenStudy (aaronq):

Post what you did and i can check it

OpenStudy (anonymous):

Then tell me if I balanced the equation right. I put 2Na + 2HF "yields" 2NaF + H2

OpenStudy (aaronq):

yes, it good

OpenStudy (anonymous):

Then I put the moles that are given 4.50 mol Na x 1mol H2/2mol Na x 2.016 H2/1mol h2, which got me 4.536

OpenStudy (aaronq):

where did the 2.016 come from?

OpenStudy (anonymous):

There are two Hydrogen so I grabbed 1.008 and added 2 of them together

OpenStudy (anonymous):

1.008 G i mean

OpenStudy (aaronq):

that's the molar mass, were not interested in finding the mass note that the answers are given in moles

OpenStudy (anonymous):

oh okay can you help me solve the equation because finding the theoretical and actual yields are very hard for me to understand.

OpenStudy (aaronq):

you pretty much had it \(\dfrac{4.5}{2}=\dfrac{moles~of~H_2}{1}\) moles of \(H_2=\dfrac{4.5}{2}=2.25~moles\) (this is the theoretical yield) Then find what 85% of this is

OpenStudy (aaronq):

Usually the percent yields are in terms of mass (grams) but you should always check the units given in the options (if there are)

OpenStudy (anonymous):

Oh my god thank you very much. How long will you be online Because I may need help in a little bit and I would love to be able to reach out to you for help.

OpenStudy (anonymous):

I may need help with a few more questions is what I meant to say

OpenStudy (aaronq):

I'm on often, i can't say how long exactly. But there are other people here who can help you as well.

OpenStudy (anonymous):

Well then is it okay if you help me with another question of finding the theoretical and percent yields?

OpenStudy (aaronq):

sure post it !

OpenStudy (anonymous):

On the same question or as a new post?